# Distilling Event Sequence Knowledge From Large Language Models

### v1

@@ 1 Introduction
Building probabilistic models from event sequence data has numerous practical applications across different domains when plentiful high-quality event data is available. For example, in Finance, event models can be used to predict stock market trends and make informed investment decisions. In Healthcare, event models can help identify patterns and correlations in patient data to improve diagnoses and treatment plans. In the field of Cybersecurity, these models can be used to detect and prevent potential cyber attacks by analyzing the sequence of events leading up to a breach. A common characteristic of the data in these domains is that sequences are clearly associated with an entity (e.g., a company, a person, or a device). There are however other domains where such a clean association between events and entities may not be possible. One such application is news event analysis Radinsky et al. (2012); Cekinel and Karagoz (2022); Du et al. (2022); Hassanzadeh et al. (2022). While various news sources record and describe newsworthy events, it is often not possible to automatically put together coherent sequences of events, because each event may involve multiple topics and actors, and many correlated and unrelated events may be occurring simultaneously or in close proximity
@@
Prior work has addressed this challenge by devising automated mechanisms for extracting narratives Norambuena et al. (2023); Santana et al. (2023), topic detection and tracking Allan (2012), and timeline summarization Gholipour Ghalandari and Ifrim (2020). While these different categories of solutions have been successful in a range of applications, the outcome is inherently noisy and not in the form of structured event sequences useful for the construction of event models. Prior work has shown little success in automatically turning narratives into structured sequences. In this paper, we explore a novel mechanism for creating structured event sequences in such domains, using generative language models.
Figure 1: Examples of frequent patterns mined from LLM-generated event sequence collections with their support (top); and some results from a simple probabilistic event model learned over the sequences that discovers influencing sets of event concepts and estimates probabilities based on historical occurrences (bottom).
@@
Large Language Models Brown et al. (2020); Raffel et al. (2020); Wei et al. (2022) have recently become the dominant paradigm in a range of natural language processing (NLP) tasks Chung et al. (2022) and often beat traditional approaches on a number of challenging tasks, including complex arithmetic reasoning Imani et al. (2023) and open-domain question answering Kamalloo et al. (2023). In this paper, our goal is to examine the capability of LLMs to generate structured event sequences useful for event analysis. Our aim is to use LLMs directly for generation of event sequences, as opposed to improving prior extraction methods that would rely on high-quality and comprehensive sources describing detailed event sequences for a wide variety of events. Our hypothesis is that LLMs trained on large corpora have already gathered the required knowledge of plausible event sequences and therefore can be suitably guided to produce diverse and high-quality event sequences. To effectively distill this knowledge, we use event-related concepts in Wikidata Vrandečić and Krötzsch (2014), a comprehensive general-domain knowledge graph, to guide the sequence generation. This can be viewed as a novel mechanism for knowledge-guided text generation Yu et al. (2022). We then use these generated sequences for pattern mining and learning probabilistic event models, as a way of structuring the underlying knowledge. Figure 1 shows examples from our experiments of patterns mined from an LLM-generated event sequence collection, as a well as a simple model learned from the collection.
@@
Upon publication, we will make our code and generated sequence data publicly available (as a separate dataset) for future research.
2 Methods
We split the task of distilling event sequence knowledge from LLMs into two sub-tasks: (i) generation of new and unique event sequences given an event trigger type; and (ii) mining high-support sequence patterns not already derivable from the knowledge graph, and learning probabilistic models to discover complex event sequence rules. Figure 2 describes our overall framework for carrying out these two objectives. We frame these two subtasks as instances of casual event prediction and event sequence mining, respectively.
Figure 2: An overview of our event sequence generation framework and an example portraying a potential use case. We show that by (1) starting with a sparse knowledge graph such as Wikidata, we can generate targeted event sequences. Owing to the inherent sparsity in the underlying KG, we can (2) use LLMs to carry out a portion of the evaluation (i.e. precision) to select an optimal model. On this new generated sequence dataset, we then (3) apply classical pattern-mining algorithms (i.e. GSP) to identify potentially interesting has_normal-_\__cause and has_normal-_\__effect event sequence chains, and (4) learn summary Markov models (SuMMs) to identify potential influencing events (i.e. parent event types) for particular event types of interest; these are both illustrations of extracting complex structured knowledge from the generated sequences.
In real-world knowledge graphs like Wikidata, such events are represented by a text-based identifier. For example, each event type in Figure 2 corresponds to a high-level newsworthy concept in Wikidata. More importantly, these event concepts are not associated with any relevant time-stamps but often do have causal relations (e.g. has_cause or has_effect) – either directly (earthquake causes tsunami) or through their instances (an earthquake event caused a tsunami event).
2.1 LLMs as Event Sequence Generators
@@
We first analyze the ability of LLMs to model event prediction as a conditional generation task under zero and iterative few-shot settings. Our targets are linearized sequences of event concepts. We begin by prompting a large language model (with in-context exemplars) with an event trigger $y_{1}$ to generate the next concept from a defined set of labels, and repeat this process until we get a sequence of desired length. Formally, given an event trigger $\mathcal{T}$, we model the probability of generating linearized string $y$ of length $T$ containing $N$ unique event concepts that follow $\mathcal{T}$ in sequence:
$p_{\text{LM}}(y|\mathcal{T})=\prod_{t=2}^{T}p(y_{t}|\mathcal{T},y_{<t-1})$
@@
This is the standard conditional language modeling objective. We try multiple prompting techniques and qualitatively observe optimal results with an iterative in-context few-shot prompting strategy (Figure 3). Specifically, we start with a set of six randomly selected examples of the form – “What usually follows event $X$?”. Based on the model output ($Y$) from a pre-specified vocabulary, we append this same example to the original prompt in conjunction (i.e. “What usually follows event $X$ and $Y$?”) with ICL examples of the same form. As shown in Figure 3, our iterative technique serves dual purposes: (i) it leverages in-context learning; and (ii) eliminates the need for implementing complex resolvers to post-process model outputs. We repeat this process until a sequence of a desired minimum length $m$ is achieved (in our experiments, $m=3$) or we’ve exhausted a maximum number of tries ($k=10$ in our experiments) to generate an in-domain event type.
@@
Figure 3: Illustration of our approach to elicit has_normal-_\__effect and has_normal-_\__cause events given a label of interest. Use of instructional in-context exemplars substantially reduces the need for post-processing LLM output in addition to constraining the output label space.
2.2 Event Concepts
@@
We curate a new dataset of high-level event concepts (classes) from Wikidata that represent newsworthy events. To do so, we query Wikidata for event concepts that have links to Wikinews articles and are instances of classes that are a subclass of the occurrence class, i.e., indicating they are newsworthy event classes. We gathered $50$ top-level classes for these event concepts, each having multiple labels (e.g. conflict $\rightarrow$ conflict (psychological), dispute, disagreement, etc.). This yielded a total of $202$ unique event labels for the $50$ top-level classes. Most of these event concepts have some causal relations (i.e. has_cause or has_effect). We use these relation pairs as in-context exemplars to create our prompts. Full length prompts used in all our experiments are provided in Appendix LABEL:sec_apx:event_seq_prompts.
Algorithm 1 Generalized Sequential Pattern Algorithm
Obtain sequences $(<x>)$ as $len-1$ candidates
$F_{1}=$ set of $len-1$ sequential patterns
$k=1$
[…]
     Generate $C_{k+1}$triangleright $len-(k+1)$ candidates from $F_{k}$
     if $C_{k+1}\neq\emptyset$  then
         Scan all sequences
         Identify $F_{k+1}$triangleright $len-(k+1)$ sequence patterns
     end if
     $k=k+1$
end while
2.3 Mining New Patterns
In order to derive new, unseen relationships (has_cause or has_effect) between the extracted event classes, we use a large pretrained language model to generate new event sequences and then use classical frequent itemset mining algorithms like GSP Srikant and Agrawal (1996) and SPADE Zaki (2004) to derive new high support patterns. This aspect of the overall workflow is illustrated in the top right section of Figure 2, where rules are mined from event sequences.
@@ 2.3 Mining New Patterns
To generate new event sequences in a zero-shot setting, we start with an event-trigger (e.g. a concept like workplace accident), create a prompt with instructions describing desired relationships and a few in-context exemplars (ICL prompt), and constrained the model output to the original $50$ event concept labels to generate the next event in the sequence. (The full text of the prompt is provided in the supplementary material.) We append the model output to another ICL prompt with conjunctive event examples (i.e. questions of the form “What happens after X, Y, and Z”). We repeat this process until we reach a pre-defined maximum sequence length (in this case, $10$) or until the model fails to generate an in-vocabulary response in $k$ maximum attempts (in this case, $k=3$). We test several ablations varying prompt length, example types, and example selection. We provide results and other details from these ablations in the supplementary material (appendix). We repeat this sequence generation procedure on all $202$ event concepts, generating $2,276$ event sequences with an average length of $5.7$ labels per sequence.
The classical pattern mining algorithms (i.e. GSP and SPADE) we use to mine for new event patterns are both a priori-based approaches (Algorithm 1). Under both methods, given two sequences of the same event concept class
$\alpha=<a_{1},a_{2},..,a_{n}>\text{ and }\beta=<b_{1},b_{2},..,b_{n}>$
$\alpha$ is a subsequence of $\beta$ denoted as $\alpha\subseteq\beta$ iff there exists a set of values $1\leq j_{1}<j_{2}<...<j_{n}\leq m$ such that $a_{1}\subseteq b_{j1},a_{2}\subseteq b_{j2},...,a_{n}\subseteq b_{jn},$; then $\beta$ is a supersequence of $\alpha$. Given multiple such sequences and a support threshold, the task here is to find a set of frequent event subsequences.
2.4 Identifying Influencing Sets through Summary Markov Models
Learning about potential influencing events that lead to a given event type in a large set of event sequences is an important aspect of analyzing multivariate event sequences, i.e. events without time stamps, like the ones generated by our models. Classic $k_{th}$ order Markov chains capture these dynamics by modeling the probability of observing a particular event type given the preceding $k$ events in-sequence. The recent family of summary Markov models (SuMMs) Bhattacharjya et al. (2022) generalize other well known Markov models for event sequences by leveraging a function that summarizes historical event occurrences, and identifying the subset of event types that affect the probability of occurrence of event types of interest; this forms the influencing set.
We use the LLM-generated sequences to learn two types of SuMMs: binary SuMMs (BSuMMs) and ordinal SuMMs (OSuMMs). Briefly, in BSuMMs, it is only the presence or absence of an event in the relevant history that has an effect on the occurrence of other events, while in OSuMMs the order of the events is also taken into account. We refer the readers to Sections 3.3 and 3.4 in Bhattacharjya et al. (2022) for formal definitions and methods for learning SuMMs over event sequence collections.
Figure 1 shows examples of learned influencing sets using BSuMMs for refugee crisis and mass migration events. Two events mass migration and famine are identified as influencing events for refugee crisis. The model indicates that the occurrence of both mass migration and famine together has a $0.91$ probability of resulting in refugee crisis as a part of an event sequence. On the other hand the occurrence of mass migration in the absence of famine has a $0.31$ probability of resulting in refugee crisis. search strategy to efficiently discover the minimal influencing sets. We view the discovery of influencing sets from LLM-generated data as a mechanism of distilling complex symbolic knowledge from the output of LLMs.
3 Experimental Settings
We performed all our model inference related experiments on two NVIDIA V100 GPUs. We used the Huggingface library v4.26.14.26.14.26.14.26.1 Wolf et al. (2020) and publicly available model checkpoints.111huggingface.co/docs/transformers/model_doc/flan-t5 Classical pattern mining algorithms (GSP, SPADE) were implemented in Python through the use of spmf222pypi.org/project/spmf/ library v1.41.41.41.4.
For generation we use in-context learning under zero shot settings for all LLMs, with top-k sampling ($k=50$, in conjunction with top-p, with $p=0.95$).
To identify influencing events from Summary Markov Models (SuMMs), we use implementations from Bhattacharjya et al. (2022) and split the dataset (generated from LLMs) into train/dev/test sets ($70\%/15\%/15\%$) to generate results reported in Table 2. BSuMMs and OSuMMs were learned with hyperparameters of $\alpha=0.1$, $\gamma=0.5$ and a look-back ($\kappa$) window of $4$.
4 Evaluation
@@
Open ended text generation in a task like event sequencing, with extremely sparse reference data, poses unique challenges to the evaluation of model outputs Wadhwa et al. (2023). The traditional scheme to evaluate discrete model outputs has been to calculate precision and recall for the generated outputs against a predefined reference test set. However, under open-world settings and particularly when a sparse KG like Wikidata is used as reference data, a missing causal relation between two event classes in a sequence may very well be a valid relation. Therefore, it is not possible to automatically derive an accurate measure of precision and recall purely using Wikidata as reference data.We take a multi-pronged evaluation approach to assess the quality and usefulness of the generated event sequences across multiple tasks.
4.1 Evaluation of Recall
Despite the sparsity of causal relations in Wikidata, one can still reasonably estimate recall through pairwise comparisons of generated event classes to the existing causal relations in Wikidata. We build a reference set of causal relations in Wikidata by curating a list of all the pairs of event concepts that have any of the several causal relations333www.wikidata.org/wiki/Wikidata:List_of_properties/causality in Wikidata in any direction, including has_cause (P828), has_immediate_cause (P1478), has_contributing_factor (P1479), and has_effect(P1542). Our results show that this is an effective mechanism of comparing recall across methods, as larger models that are expected to have better accuracy yield higher recall scores.
4.2 Evaluation of Precision
To overcome sparsity in reference data, prior work has generally relied on human evaluations Chiang and Lee (2023)for estimating precision, which entails looking at pairs of events and asking annotators whether or not the events in question have a causal relationship. Such a process for potentially thousands of event pairs (like in our case) can be very cost prohibitive. Furthermore, recent research Zhao et al. (2022); He et al. (2023) indicates that pretrained language models themselves might outperform lay human annotators such as those found on Amazon Mechanical Turk. For instance, He et al. (2023) demonstrated that labeling data under a few-shot chain-of-thought prompt (“explain-then-annotate” setting) surpasses crowdworker annotations on relevance assessments. Following these results, we use LLMs for evaluating precision and for selecting the most optimal model.
We propose evaluating precision for model selection as a binary classification task. Given two events $e_{1},e_{2}$, an evaluator model must evaluate whether $e_{1}$ reasonably leads to $e_{2}$ under a has_cause or has_effect relationship. To do this, we start with a large instruction-tuned model (in our case, Flan-T5-XXL ($11$B) Chung et al. (2022)) and adopt instructional in-context few shot prompting to classify whether or not $e_{1}$ reasonably leads to $e_{2}$, and for the model to provide a justification for its results. While we hypothesize that such an approach may yield noisy results, in a preliminary manual assessment this approach seems to yield reasonable outputs that are sufficient for evaluating correctness of model outputs (i.e. precision). We include a sample of model evaluated outputs in supplementary material, and select few examples in Appendix LABEL:sec_apx:examples.
Precision Evaluator Model (P) R F-1
Flan-T5-Large Flan-T5-XL Flan-T5-XXL
Flan-T5-Large ($783$M) Chung et al. (2022) 0.73 0.72 0.72 0.47 0.57
Flan-T5-XL (3B) 0.75 0.75 0.78 0.49 0.60
Open-Research LLaMA ($3$B) Touvron et al. (2023) 0.70 0.69 0.72 0.45 0.55
Flan-T5-XXL ($11$B) 0.79 0.79 0.81 0.54 0.65
[…]

### v3

@@ Abstract
Keywords:
Knowledge Graphs, Large Language Models, Knowledge Distillation
@@ 1 Introduction
Building probabilistic models from event sequence data has numerous practical applications across different domains when plentiful high-quality event data is available. For example, in Finance, event models can be used to predict stock market trends and make informed investment decisions. In Healthcare, event models can help identify patterns and correlations in patient data to improve diagnoses and treatment plans. In the field of Cybersecurity, these models can be used to detect and prevent potential cyber attacks by analyzing the sequence of events leading up to a breach. A common characteristic of the data in these domains is that sequences are clearly associated with an entity (e.g., a company, a person, or a device). There are however other domains where such a clean association between events and entities may not be possible. One such application is news event analysis [8, 13, 18, 30]. While various news sources record and describe newsworthy events, it is often not possible to automatically put together coherent sequences of events, because each event may involve multiple topics and actors, and many correlated and unrelated events may be occurring simultaneously or in close proximity
@@
Prior work has addressed this challenge by devising automated mechanisms for extracting narratives [29, 33], topic detection and tracking [2], and timeline summarization [16]. While these different categories of solutions have been successful in a range of applications, the outcome is inherently noisy and not in the form of structured event sequences useful for the construction of event models. Prior work has shown little success in automatically turning narratives into structured sequences. In this paper, we explore a novel mechanism for creating structured event sequences in such domains, using generative language models.
@@
Large Language Models [7, 31, 38] have recently become the dominant paradigm in a range of natural language processing (NLP) tasks [11] and often beat traditional approaches on a number of challenging tasks, including complex arithmetic reasoning [22] and open-domain question answering [23]. In this paper, our goal is to examine the capability of LLMs to generate structured event sequences useful for event analysis. Our aim is to use LLMs directly for generation of event sequences, as opposed to improving prior extraction methods that would rely on high-quality and comprehensive sources describing detailed event sequences for a wide variety of events. Our hypothesis is that LLMs trained on large corpora have already gathered the required knowledge of plausible event sequences and therefore can be suitably guided to produce diverse and high-quality event sequences. To effectively distill this knowledge, we use event-related concepts in Wikidata [36], a comprehensive general-domain knowledge graph, to guide the sequence generation. This can be viewed as a novel mechanism for knowledge-guided text generation [42] and symbolic knowledge distillation [39]. We then use these generated sequences for pattern mining and learning probabilistic event models, as a way of further structuring the underlying knowledge. Figure 1 shows our overall framework along with examples from our experiments of patterns mined from an LLM-generated event sequence collection, as a well as a simple model learned from the collection.
Figure 1: An overview of our framework for distilling event sequence knowledge from LLMs, along with examples portraying potential use cases. We show that by (1) starting with a sparse knowledge graph such as Wikidata, we can generate targeted event sequences. Owing to the inherent sparsity in the underlying KG, we can (2) use LLMs to carry out a portion of the evaluation (i.e. precision) to select an optimal model. On this new generated sequence dataset, we then (3) apply classical pattern-mining algorithms (e.g., GSP) to identify potentially interesting has__\__cause and has__\__effect event sequence chains, and (4) learn summary Markov models (SuMMs) to identify potential influencing events for particular event types of interest; these are both illustrations of extracting complex structured knowledge from the generated sequences.
@@
Our code and generated sequence data are included in the supplementary material, and will be made publicly available for future research.
2 Related Work
News Event Analysis
The primary applications of our work are around news event analysis and forecasting. Liang [45] presents a taxonomy of different flavors of event prediction in the literature. Our target event prediction applications fall under the “Semantic Prediction" category, with time and location details not being of interest, and the primary goal being the prediction of “event profiles" such as event types. Seminal work in this area is the work of Radinsky et al. [30] where causal relations between past events are extracted from text and then a knowledge graph is utilized to generalize the extracted relations in order to make predictions. More recent work has explored the use of graph sequence mining over a graph structure representation of events extracted from text [8], with graph mining used as a mechanism of extracting useful relations from large and noisy outputs of extraction. The application of building event sequence models over such outputs has not been explored. Our paper explores an alternative approach of event sequence generation that is capable of generating longer and potentially higher-quality sequences.
Event Sequence Extraction from Text
There is a wealth of literature on different methods of extracting sequences from textual corpora. Norambuena et al. [29] present a comprehensive survey of automated methods of narrative extraction. Narratives contain several elements including events, participants (actors/protagonists), time, and space. The task of event detection itself is a highly challenging task and the topic of extensive research [9, 41, 24], which has its root in the Topic Detection and Tracking (TDT) task [33]. This line of work started out as a DARPA-sponsored initiative with the same name [3]. Another closely related task is news timeline summarization [16]. While pattern mining algorithms have been applied to the output of such extractions, e.g. for creation of “domain templates" [14], we are not aware of any attempts to use the extractions to construct event models for analysis or prediction. This is mainly due to the fact that the output of such methods often result in very short and noisy sequences, not suitable for the majority of event sequence models.
Sequential Pattern Mining and Event Sequence Models
Sequential pattern mining and related approaches have been the subject of extensive research [27, 26, 28, 15]. These algorithms take in a set of sequences (or “sequential records") with a set of unique events (or “items") and often a “minimum support" threshold, and return as output a ranked set of all frequent sequences in a given sequence collection (or database) meeting the minimum support threshold. There is also a long line of work on statistical and probabilistic models for analysis of different kinds of event sequences. Our focus in this paper is on multivariate event sequences, i.e., sequences of various event types without timestamps. Markov models for sequences [32, 5] and long short-term memory (LSTM) [20] models are examples of prediction models. We leverage a more recent family of models – Summary Markov models [6] – for some of the experiments in this paper that involve analyzing generated event sequences.
3 Knowledge-Guided Event Sequence Generation
@@
Through utilizing LLMs, we model event prediction as a conditional generation task under zero and iterative few-shot settings. Our targets are linearized sequences of event concepts. We begin by prompting a large language model (with in-context exemplars) with an event trigger $y_{1}$ to generate the next concept from a defined set of labels, and repeat this process until we get a sequence of desired length. Formally, given an event trigger $\mathcal{T}$, we model the probability of generating linearized string $y$ of length $T$ containing $N$ unique event concepts that follow $\mathcal{T}$ in sequence:
$p_{\text{LM}}(y|\mathcal{T})=\prod_{t=2}^{T}p(y_{t}|\mathcal{T},y_{<t-1})$ (1)
@@
This is the standard conditional language modeling objective. We try multiple prompting techniques and qualitatively observe optimal results with an iterative in-context few-shot prompting strategy (Figure 2). Specifically, we start with a set of six randomly selected examples of the form – “What usually follows event $X$?”. This approach follows the incremental prompting procedure from [25]. Based on the model output ($Y$) from a pre-specified vocabulary, we append this same example to the original prompt in conjunction (i.e. “What usually follows event $X$ and $Y$?”) with ICL examples of the same form. As shown in Figure 2, our iterative technique serves dual purposes: (i) it leverages in-context learning; and (ii) eliminates the need for implementing complex resolvers to post-process model outputs. We repeat this process until a sequence of a desired minimum length $m$ is achieved (in our experiments, $m=3$) or we’ve exhausted a maximum number of tries ($k=10$ in our experiments) to generate an in-domain event type.
@@
Figure 2: Illustration of our approach to elicit event sequence knowledge given a label of interest. Use of instructional in-context exemplars substantially reduces the need for post-processing LLM output in addition to constraining the output label space.
Identifying Event Concepts
@@
With incremental prompting we curate a new dataset of high-level event concepts (classes) from Wikidata that represent newsworthy events. To do so, we query Wikidata for event concepts that have links to Wikinews articles and are instances of classes that are a subclass of the occurrence class, i.e., indicating they are newsworthy event classes. We gathered $50$ top-level classes for these event concepts, each having multiple labels (e.g. conflict $\rightarrow$ conflict (psychological), dispute, disagreement, etc.). This yielded a total of $202$ unique event labels for the $50$ top-level classes. Most of these event concepts have some causal relations (i.e. has_cause or has_effect). We use these relation pairs as in-context exemplars to create our prompts. We then generate event sequences through iterative in-context prompting (Figure 2). Full length prompts used in all our experiments are provided in the Appendix.
@@
To generate new event sequences in a zero-shot setting, we start with an event trigger (e.g. a concept like workplace accident), create a prompt with instructions describing desired relationships and a few in-context exemplars (ICL prompt), and constrained the model output to the original $50$ event concept labels to generate the next event in the sequence (full text of the prompt is provided in the supplementary material). We append the model output to another ICL prompt with conjunctive event examples (i.e. questions of the form “What happens after X, Y, and Z”). We repeat this process until we reach a pre-defined maximum sequence length (in this case, $10$) or until the model fails to generate an in-vocabulary response in $k$ maximum attempts (in this case, $k=3$). In this process, we test the following two ablations –
Number of exemplars
We varied the number of in-context exemplars between 1-12. This is in addition to the zero-shot experiments described in the main paper. We evaluated recall (i.e. proportion of references captured by the resulting output sequences) for all generated outputs in an automated way through matching lexical alignment with Wikidata references. We observed none to marginal improvements by varying the number of exemplars beyond 3 in the initial trigger prompt, and beyond 5 in the second iterative prompt.
Selection of specific exemplars
Selecting in-context examples is an incredibly noisy process [44]. We started with a static set of randomly selected examples, however owing to Wikidata’s inherent label imbalance (political and economic events dominate newsworthy concepts), this led to erratic results, i.e. high recall on similar concepts but not otherwise.
We then tested a dynamic selection method where every instance of the prompt would contain examples similar to target__\__label. To achieve this, we used BERT embeddings [12] to retrieve (using cosine distance) examples from the reference set. This approach however led to a lower macro-recall and, upon manual inspection of outputs, we observed a high degree of redundancy, i.e. generated outputs were copied from the ICL exemplars.
To solve for these issues, we reverted to the static prompt examples but manually selected a mix of event topics to be included in the prompt. Our current selection of the prompt yields higher macro-recall than both of the aforementioned techniques tested. While we believe there may be better techniques to select ideal candidates for ICL exemplars [4, 1] and understand how their compositionality affects specific outputs, we consider such an analysis to be beyond the scope of our work.
We repeat this sequence generation procedure on all $202$ event concepts, generating $2,276$ event sequences with an average length of $5.7$ labels per sequence.
4 Assessing the Quality of LLM-Generated Event Sequences
@@
Open-ended text generation in a task like event sequencing, with extremely sparse reference data, poses unique challenges to the evaluation of model outputs [37]. The traditional scheme to evaluate discrete model outputs has been to calculate precision and recall for the generated outputs against a predefined reference test set. However, under open-world settings and particularly when a sparse KG like Wikidata is used as reference data, a missing causal relation between two event classes in a sequence may very well be a valid relation. Therefore, it is not possible to automatically derive an accurate measure of precision and recall purely using Wikidata as reference data. We take a multi-pronged evaluation approach to assess the quality and usefulness of the generated event sequences across multiple tasks. This section presents our approach in evaluating the quality of generated event sequences, along with the results of this evaluation. An evaluation of the usefulness of the generated event sequences is presented in Section 5.